# Self-Supervised Enhancement of Forward-Looking Sonar Images: Bridging Cross-Modal Degradation Gaps through Feature Space Transformation and Multi-Frame Fusion

### v2

@@ IV Experiments
In this section, we demonstrate that our method outperforms existing approaches [15, 25] on a real-world forward-looking sonar dataset containing three target types. Our experiments, conducted on a 24GB NVIDIA GeForce GTX 4090, show that the accurate pose information in the sonar images eliminates the need for additional registration before multi-frame fusion, allowing 16 images to be fused in just 0.3 seconds during inference.
@@ IV-A Forward-Looking Sonar Datasets
We built a standard data collection pool for forward-looking sonar (see Figure 5) using a C900-II unit operating at 2.7 MHz, with a horizontal beamwidth of 50° and a vertical beamwidth of 15°. The sonar is fixed 1.5 m from the guide rail cantilever and tilted 15° downward relative to the rotating platform. During acquisition, the system performs a 360° rotation around the target, with the guide rail rotation angle recorded continuously. Combining this rotation data with the initial sonar pose enables calculation of the sensor’s pose at each instance, eliminating the need for additional image registration prior to multi-frame fusion.
The dataset comprises multi-frame forward-looking sonar images capturing three target types from various materials and viewpoints: rubber tires (exhibiting severe acoustic absorption), metal torpedo models, and glass-reinforced plastic (GRP) conical frustum. Each target class contains 3,200 images grouped into sets of 16 frames. We randomly selected 180 groups for training and 10 groups each for validation and testing. All images have been converted from polar to Cartesian coordinates at a resolution of 376 × 376.
@@ IV-B Evaluation Metrics for Enhancement Performance
In underwater environments, high-quality reference images are scarce, making common metrics for optical image enhancement, such as PSNR, unsuitable. We adopt methods from [24, 25] to evaluate the denoising performance and the preservation of high-frequency details in sonar image enhancement. Standard deviation (STD) is used to assess image smoothness—lower STD indicates stronger noise suppression—while average gradient (AG) measures high-frequency detail, with higher AG reflecting richer detail.
@@ IV-C Quantitative and Visual Evaluation of Enhancement Results
For the torpedo model, the STD is reduced from 17.02 to 13.14, surpassing wavelet-based (21.59) and HighRes-based (14.58) methods. AG improves from 4.93 to 5.92, indicating sharper edges and clearer target contours.
@@
Fusion Method Original Wavelet [15] Highres [25] Proposed
@@
Fusion Method Original Wavelet [15] Highres [25] Proposed
@@
Fusion Method Original Wavelet [15] Highres [25] Proposed
@@ IV-D Ablation Study
To further assess the contribution of the proposed feature space transformation module, we conducted ablation studies within our multi-frame self-supervised enhancement framework, incorporating several widely used feature operators in sonar image processing. These include Histogram of Oriented Gradients (HOG), Canny edge detection, Gray-Level Co-occurrence Matrix (GRE), Haar wavelet features (HAAR), and Wavelet Scattering Transform (WST).
@@
Table II reports the quantitative results using standard deviation (STD) and average gradient (AG). The results highlight notable performance differences. Canny achieved the lowest STD, indicating effective noise suppression. However, its low AG suggests over-smoothing and the loss of critical high-frequency edge details.
In contrast, WST delivered the most balanced performance, combining strong noise reduction with superior detail preservation. This is largely due to its multi-scale analysis capability, which enables robust extraction and retention of structural features, particularly in low-quality forward-looking sonar images.
Figure 7 further visualizes the enhancement results of Canny and WST. The WST-based output exhibits higher overall intensity, clearer boundaries, and richer detail, while the Canny-based result shows evident blurring and loss of edge information. Such high-frequency detail and contrast enhancement are essential for accurate underwater target detection.
In conclusion, the ablation study demonstrates that WST provides the optimal trade-off between denoising and detail retention, making it the most effective operator for bridging the cross-modal degradation gap. As such, WST is selected as the default feature transformation module in our final approach.
@@ V Conclusion
This paper presents a self-supervised enhancement method for forward-looking sonar images, combining feature space transformation with multi-frame fusion. The proposed transformation module effectively bridges the cross-modal degradation gap between sonar and other domains (e.g., multispectral remote sensing), enabling a practical transfer learning strategy. A self-supervised fusion network leverages the high frame rate of sonar systems to integrate complementary frame information, naturally suppressing noise while preserving high-frequency details. Experiments on three real-world, heavily degraded sonar datasets show that the method outperforms existing approaches in both quantitative metrics and visual quality. With an inference time of 0.3 seconds for 16 frames, the method is also efficient and well-suited for deployment on underwater vehicles to enhance target detection.

### v3

@@ IV Experiments
Most existing sonar image enhancement methods rely on supervised learning with simulated data, which often fails to generalize to real-world scenarios. In contrast, we compare against two recent state-of-the-art reference-free techniques—SEGSID [24] and Highres-based method  [25]—both of which do not require high-quality reference images. We evaluate all methods on a self-collected forward-looking sonar dataset featuring three target types. Our experiments, conducted on a 24 GB NVIDIA GeForce GTX 4090, leverage accurate pose information to avoid any additional registration before multi-frame fusion. By closing the gap between sonar and remote-sensing imagery, our Deformable WST Feature Bridge enables the direct fine-tuning of hyperspectral remote-sensing pretrained weights from [46], accelerating convergence and improving enhancement quality on real sonar data. Coupled with accurate pose metadata, this adaptation allows the fusion of 16 frames in just 0.3 s at inference.
@@ IV-A Forward-Looking Sonar Datasets
We built a standard data collection pool for forward-looking sonar (see Figure 4) using a C900-II unit operating at 2.7 MHz, with a horizontal beamwidth of 50° and a vertical beamwidth of 15°. The sonar is fixed 1.5 m from the guide rail cantilever and tilted 15° downward relative to the rotating platform. During acquisition, the system performs a 360° rotation around the target, with the guide rail rotation angle recorded continuously. Combining this rotation data with the initial sonar pose enables calculation of the sensor’s pose at each instance, eliminating the need for additional image registration prior to multi-frame fusion. Each target type comprises 3,200 sonar images grouped into sequences of 16 frames each. We randomly selected 180 groups for training, and reserved 10 groups each for validation and testing. All images were converted from polar to Cartesian coordinates at a resolution of 376 × 376 pixels. Moreover, instead of training our fusion network entirely from scratch on sonar data, we initialize it with the pretrained weights provided by [25], thereby leveraging existing remote‐sensing models for faster convergence and improved generalization.
Despite being acquired in a controlled environment, our sonar images exhibit severe degradation reflective of realistic underwater scenarios, including multipath scattering, acoustic absorption, and intrinsic speckle noise. The dataset features three targets representing typical degradation conditions: rubber tires with severe acoustic absorption causing low contrast and blurred edges; metal torpedo models exhibiting complex multipath reflections and intense speckle noise; and a glass-reinforced plastic (GRP) conical frustum with moderate acoustic scattering characteristics commonly encountered in underwater missions. A circular trajectory ensuring full 360° coverage around each target was adopted, providing diverse angular views and accurate pose estimation without additional image registration, essential for evaluating multi-frame fusion robustness. Thus, our dataset effectively represents real-world sonar imaging challenges, validating the robustness and applicability of our enhancement method.
@@ IV-B Evaluation Metrics for Enhancement Performance
In underwater environments, high‐quality reference images are scarce, making common metrics for optical image enhancement (e.g., PSNR, SSIM) impractical. Instead, we follow established reference-free sonar enhancement works [24, 25] and employ two no‐reference metrics that capture complementary aspects of image quality:Standard deviation (STD) is used to assess image smoothness—lower STD indicates stronger noise suppression—while average gradient (AG) measures high-frequency detail, with higher AG reflecting richer detail.
@@ IV-C Quantitative and Visual Evaluation of Enhancement Results
For the torpedo model, the STD is reduced from 17.02 to 13.14, surpassing SEGSID (21.59) and HighRes-based (14.58) methods. AG improves from 4.93 to 5.92, indicating sharper edges and clearer target contours.
@@
Fusion Method Original SEGSID [24] Highres [25] Proposed
@@
Fusion Method Original SEGSID [24] Highres [25] Proposed
@@
Fusion Method Original SEGSID [24] Highres [25] Proposed
@@
Figure 6: Visualization of handcrafted features. WST is selected for its balanced performance in denoising and detail retention.
@@ IV-D Ablation Study
Figure 6 shows the handcrafted feature operators (HOG, Canny, GRE, HAAR, WST). We perform ablation experiments within our multi-frame reference-free enhancement framework to demonstrate that WST achieves the best balance between noise suppression and detail preservation, justifying its selection as the primary feature operator.
@@
Table II reports quantitative results in terms of standard deviation (STD, lower is better for noise suppression) and average gradient (AG, higher is better for detail preservation). While Canny achieves the lowest STD—indicating strong noise removal—its AG is drastically lower than other methods, reflecting significant loss of edge information. In contrast, WST attains a near-optimal STD reduction and the highest AG, striking the best balance between denoising and detail retention. Figure 7 illustrates this: WST enhances intensity and edge clarity, whereas Canny’s output, despite being smoother, suffers from blurred boundaries. Hence, although Canny suppresses noise more aggressively, WST’s balanced performance makes it the preferred choice for our feature transformation module.
@@ V Conclusion
We propose a reference-free framework for forward‐looking sonar enhancement that integrates a Deformable WST Feature Bridge with an end‐to‐end multi‐frame fusion network. The Deformable WST bridge uses learnable wavelet scale and orientation offsets to align sonar features with pretrained remote‐sensing distributions, closing the cross‐modal gap. The fusion network then aggregates complementary frames to suppress speckle noise and boost target brightness—without any external ground truth. On three self‐collected datasets (rubber tire, metal torpedo, GRP frustum), our method outperforms all reference‐free baselines in STD, AG, and visual quality. By narrowing the degradation gap, it also allows direct fine‐tuning of remote‐sensing pretrained weights for faster convergence and stronger generalization. Achieving 16‐frame processing in just 0.3 s on a single GPU, it is well suited for real‐time underwater deployment. Future work will validate the approach in field environments and embed it within autonomous underwater vehicles.